# Generating Hypotheses of Dynamic Causal Graphs in Neuroscience: Leveraging Generative Factor Models of Observed Time Series

### v1

@@ 4.1 Synthetic Data Generation and Analyses
We rely on synthetic data to test REDCLIFF-S’ ability to estimate known causal relationships. To create these datasets, we defined each artificial temporal system as a set of Vector Auto-Regressive (VAR) models  (Zivot & Wang, 2006), one for each factor. Our implementation allows for inter-variable edge activations between time-lags to be either linear or nonlinear; in this work, we included nonlinear activations between different time-lags. Specifically, we used the ReLU (i.e. positive indicator $1(x)_{[\mathrm{max}(x,0)]}$) and negative ReLU (i.e. negative indicator $1(x)_{[\mathrm{min}(x,0)]}$) nonlinear activations. For each experiment we selected hyperparameters such as the number of VAR models used, how many time series were present, and how many lags were in each VAR (see Appendix D). Here we set the number of lags for each VAR to $\tau=2$ time steps for visualization purposes, but any integer greater than 0 would suffice. Defining these systems also meant selecting square adjacency matrices mapping values from past to present system states for each VAR model in the system. We included self-connections in each adjacency matrix to better visualize their generated time series. The number of inter-variable connections between time series was varied across systems, with connections chosen randomly. The sizes of our Synthetic Systems were primarily designed to approximate the conditions of the TST dataset (Carlson et al., 2023) with increasing fidelity/complexity (the largest Synthetic Systems had 12 nodes, which corresponded to 12 channels in our pre-processed TST dataset).
Once we defined a system, we sampled recordings and their labels for the dataset. First, a random state vector was drawn from a uniform distribution for each VAR model in the system. This vector was recurrently fed through its corresponding VAR model for a number of ‘burn-in’ steps, after which we recorded the resulting multivariate time series for $T$ time steps; innovations were simulated throughout by multiplying draws from uniform and Gaussian distributions. Recordings obtained from each VAR model were then weighted over time by linearly interpolated weights (randomly selected between 0 and 1). These weighted recordings were added together along with a level of Gaussian noise. We augmented labels by marking with a one-hot vector which factor weight was largest at each time step, meaning that all of our Synthetic Systems experiments contain label noise.
@@
where $r_{\mathrm{graph}}$ is the ratio of true inter-variable interactions in each system state to the number of possible inter-variable interactions. In essence, $\mathfrak{C}$ is inversely related to the sparsity of relationships in the graph. We provide additional details regarding this complexity score in Appendix C and D.
@@ 4.2 Hypothesis Generation for Synthetic Systems
Figure 2 gives results from one high complexity system (6-2-2) and three moderate system configurations for which we generated five random repeats (i.e. unique systems) to train each algorithm (more results in Appendix C). Each repeat was designed to have 1040 training samples and 240 validation samples for each factor/system state in the system, with each sample recording 100 sequential time steps. We used the cMLP and cLSTM (Tank et al., 2021); DGCNN (Song et al., 2018); and dCSFA-NMF (Talbot et al., 2023; Mague et al., 2022) methods as baselines. Hyperparameters for each algorithm were chosen via grid search on a single repeat in our (low-complexity) $\mathrm{n_{c}}=12$-node, $\mathrm{n_{e}}=33$-edge (inter-variable), $\mathrm{n_{k}}=3$-factor Synthetic System dataset. Our grid search selected parameters which minimized stopping criteria (for DYNOTEARS and REDCLIFF-S) or objective functions (for all other baselines), with ties won by the most ‘parsimonious’ parameters (see Appendices D and E). The same hyperparameters were used across all the experiments reported in this section, with limited adjustments - e.g. to the number of REDCLIFF-S factors or the value of loss coefficients - made to account for certain characteristics in the target system (see Appendix E). Each algorithm was trained on a given system repeat dataset, after which we computed the average optimal f1-score (and corresponding standard error of the mean, or SEM) of each algorithm’s inter-variable causal predictions against the true causal graphs across factors and repeats for each synthetic system.
@@
Table 2: Comparing unweighted inter-variable edge detection by supervised discovery methods on the 12-11-2 Synthetic Systems (Fig. 2) - note REDLCIFF-S’ f1-scores differ from Fig. 2 which included edge weighting. Edge predictions were determined by optimal f1-score thresholds. Values report the mean (over repeats) of the median (over 10 random seeds and system factors) statistic $\pm$ SEM. ‘Upper’ and ‘Lower’ refer to upper- and lower-triangular portions of adjacency matrices, which we split to avoid cycles (see Sec. 4.1 & Appendix B.2). We also report the mean comparative placement (from 1 to 5) of each method’s mean performance (details in Appendix C.3).
@@ 4.3 D4IC Multi-State Extension of DREAM4
Hyperparameters for each algorithm were tuned to a single repeat of the D4IC MSNR dataset prior to training on all D4IC noise levels and repeats. Model selection proceeded as in Section 4.2, though we increased the weight on the cosine similarity penalty in selecting our final REDCLIFF-S model which balanced the scale of variation of the cosine similarity penalty with that of the other penalties (details in Appendix B.3). We include two cMLP baseline models (v1 and v2) with slightly different hyperparameters. Results are visualized in Figure 4, in which we see the REDCLIFF-S model improves over alternative approaches on this dataset.
@@ 4.4 Discussion
The main strength of REDCLIFF-S is that its hypotheses can be easily tested. Equations 4 and 8 are designed for this in several ways. As discussed, factor scores $\alpha_{n,k,t}$ can be given explicit behavioral designations. Factor scores can also be constrained (via Sigmoid or similar activation functions) to be non-negative, and thereby be viewed as how (physically) present the corresponding factors are in predicted signals. [^5] Our early experiments included these constraints, but we omit them here due to different models being selected for by our criteria. Thirdly, trained factors $f_{\phi_{k}}$ rely on fixed graphs explicitly showing relationships between variables. Finally, the sum over factors allows us to inspect individual contributions of each factor. Overall, REDCLIFF-S shows performance improvements across many scenarios.
@@
First, Figure 2 and Table 2 together give the core results on our Synthetic Systems Dataset experiments. Figure 2A presents average optimal f1-scores (across all samples and repeats) of REDCLIFF-S and applicable baselines. Surprisingly, the single-factor baselines attain fairly competitive f1-scores in some cases, seemingly learning an ‘average factor graph’ that may not be particularly accurate for any single factor but which performs well ‘on average’. Even so, in Figure 2B we see the average pair-wise improvement of REDCLIFF-S’ predictions is at least a standard error of the mean above zero for all baselines on all four systems, suggesting REDCLIFF-S yields improvements with statistical significance. These findings are further supported by the results shown in Table 2. Despite the fact that REDCLIFF-S was the only algorithm trained to predict global state information (all other baselines in Table 2 had direct access to the state label), REDCLIFF-S is uniquely able to attain competitive performance across all metrics used. This can be seen in that when we rank the algorithms based on their mean performance for each metric, REDCLIFF-S’ ranking (or ‘comparative placement’, to avoid confusion with ‘matrix rank’) tends to be lower than the other methods’. Taken together, our findings in Section 4.2 suggest REDCLIFF-S generates hypotheses about the presence of dynamic causal relationships that prove accurate at state-of-the-art rates, and this pattern holds true for systems with differing numbers of variables, inter-variable relationships, and factors.
@@
Results from our D4IC HSNR experiments are shown in Figure 4, where we see the REDCLIFF-S (cMLP) model attains the highest average optimal f1-score with a statistically significant margin. While the improvements over competing methods are somewhat modest in this case, a brief investigation into the factor networks of D4IC HSNR suggests that four of the five factors in each repeat would have been rated as ‘low complexity’ according to Equation 11; importantly, REDCLIFF-S tends to make the least improvement over baselines in the ‘low complexity’ regime (see Appendix C.3). In all, our D4IC experimental results generally support those from Section 4.2, verifying that REDCLIFF-S is effective at hypothesizing dynamic causal relationships.
@@ 5 Experiments on Real-World Datasets
We tested REDCLIFF-S on real-world data using the publicly available Tail Suspension Test (TST) dataset (Carlson et al., 2023) - which we discuss here - and the Social Preference (SP) dataset described in Mague et al. (2022) (see Appendix C.5). Where possible, we combined multiple recordings within a single brain region into an averaged LFP for that region. The number of factors and other hyperparameters were chosen based on grid searches minimizing the stopping criteria for training. Appendix D covers TST preprocessing in detail and Appendix C.4 elaborates on TST model selection and additional results.
In Figure 5 we show the difference between mean normalized causal estimates for the REDCLIFF-S factors assigned to the Open Field (OF) and Home Cage (HC) behavioral paradigms in this case study. In it, we see the OF-factor is more heavily reliant than the HC-factor on information from the Nucleus Accumbens Core (Acb_Core) to predict behavior in the Prelimbic and Infralimbic cortices, and from the Anterior-Left Lateral Habenula (‘alLH_HAB’) and Medial Dorsal Hippocampal (mDHip) channels to predict activity in several regions including the Thalmus (Md_Thal). Interestingly, our model’s hypothesis that the Anterior-Left Lateral Habenula (‘alLH_HAB’) is important for predicting LFP activity across several brain regions in the more stressful OF paradigm appears to be validated by recent research implicating the Lateral Habenula in stress responses of mice (Tan et al., 2024). Prior research also identified a brain-wide ‘anxiety network’ involving the Nucleus Accumbens Shell and the Prelimbic and Infralimbic cortices (Talbot et al., 2023), which our REDCLIFF-S model appears to express in part as pathways from the Nucleus Accumbens Core to the Prefrontal Cortex and from there to the Infralimbic Cortex and then to the Nucleus Accumbens Shell.
@@ 6 Conclusion
We present the novel REDCLIFF-S hypothesis generation algorithm for temporal systems featuring dynamic causal interactions. By learning to combine nonlinear factors using weights conditioned on signal history, REDCLIFF-S attains state-of-the-art performance in estimating multiple causal graphs simultaneously in various settings, and in a way conducive to follow-up scientific inquiry. Experiments on real data suggest REDCLIFF-S can provide scientific value to computational neuroscience and other fields.
@@ Acknowledgements
We would like to thank our colleagues at the Collective for Psychiatric Neuro-Engineering (CPNE) for supporting this work. In particular, conversations with Noah Lanier, Aaron Fleming, Kathryn Walder, and Hannah Soliman proved especially helpful in clarifying and communicating our ideas. Conversations with Natalie Brown were also useful for clarifying key ideas. We would also like to thank the reviewers for providing their insight and helpful feedback.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2139754.
@@ A.1 Identifiability Analysis of Systems Featuring Nonlinear, Dynamic Causal Graphs
In this section we provide theoretical arguments as to why the systems at the center of our work - dynamical systems featuring nonlinear, dynamic causal graphs - are generally not identifiable. At the same time, we point out that prior work has demonstrated that utilizing causal discovery methods can still benefit scientific research in real-world settings where nonlinear, dynamic causal graphs are the norm rather than the exception (Mague et al., 2022). Throughout our discussion, we follow the common definition of identifiability, in which if the likelihood of two models is the same, then their parameterizations must also be the same (Maclaren & Nicholson, 2019).
@@
There are three cases related to causal identifiability in our setting: the case where auxiliary variables are not available (the unsupervised case), that where some - but not all - auxiliary variables are available (the partially-supervised case), and the case where an auxiliary variable can be assigned to every system state (the fully supervised case). For the unsupervised and partially-supervised cases, the lack of identifiability is readily apparent in that we can swap various factors and factor weights tied to unsupervised states within our REDCLIFF(-S) architecture and arrive at functionally identical models.
@@
The fully supervised case is also not identifiable, as we illustrate via counter-example. Consider the system with $n_{\mathrm{k}}=2$ states (00 and $1$), $n_{\mathrm{c}}=2$ node ($\mathrm{A}$ and $\mathrm{B}$) where both nodes are self-connected. Assume all innovations in $\mathrm{A}$ and $\mathrm{B}$ follow a symmetric, random distribution (such as the uniform or normal distributions) in both states 00 and $1$ at every time step. Finally, assume the only difference between states 00 and $1$ is that in state 00 we have that $\mathrm{A}$ drives $\mathrm{B}$ (i.e. $\mathrm{A}\rightarrow\mathrm{B}$) via the ReLU (i.e. positive indicator $1(x)_{[\mathrm{max}(x,0)]}$) function at lag $t=2$, whereas in state $1$ we have $\mathrm{A}\rightarrow\mathrm{B}$ is the negative ReLU (i.e. negative indicator $1(x)_{[\mathrm{min}(x,0)]}$) function at lag $t=2$. In this setting the functional form of the $\mathrm{A}\rightarrow\mathrm{B}$ edge cannot be uniquely identified for either state 00 or $1$ even with an infinite number of samples. Specifically, for state 00 the negative indicator $1(x)_{[\mathrm{min}(x,0)]}$ at lag $t=2$ will be just as correlated with the observed data as the true positive indicator $1(x)_{[\mathrm{max}(x,0)]}$ at lag $t=2$ for the $\mathrm{A}\rightarrow\mathrm{B}$ edge. Similarly, in state $1$ the positive indicator $1(x)_{[\mathrm{max}(x,0)]}$ at lag $t=2$ will be just as correlated with observations as the true negative indicator $1(x)_{[\mathrm{min}(x,0)]}$ at lag $t=2$ for the $\mathrm{A}\rightarrow\mathrm{B}$ edge.
@@
In summary, we have demonstrated that systems featuring nonlinear, dynamic causal graphs are not identifiable unless one makes sweeping assumptions which preclude even fairly simple systems (such as our 2-node system with ReLU-like nonlinearities) from being considered. Despite this, prior work has shown that causal discovery methods still benefit scientific research in real-world settings where nonlinear, dynamic causal graphs are common (Mague et al., 2022).

### v2

@@ 4.1 Synthetic Data Generation and Analyses
We rely on synthetic data to test REDCLIFF-S’ ability to estimate known causal relationships. To create these datasets, we defined each artificial temporal system as a set of Vector Auto-Regressive (VAR) models  (Zivot & Wang, 2006), one for each factor. Our implementation allows for inter-variable edge activations between time-lags to be either linear or nonlinear; in this work, we included nonlinear activations between different time-lags. Specifically, we used the ReLU (that is $\mathrm{ReLU}(x)=x\mathbf{1}_{[\mathrm{max}(x,0)]}(x)$) and negative ReLU (that is $\mathrm{nReLU}(x)=x\mathbf{1}_{[\mathrm{min}(x,0)]}(x)$) nonlinear activations. For each experiment, we selected hyperparameters such as the number of VAR models used, how many time series were present, and how many lags were in each VAR (see Appendix D). Here we set the number of lags for each VAR to $\tau=2$ time steps for visualization purposes, but any integer greater than 0 would suffice. Defining these systems also meant selecting square adjacency matrices mapping values from past to present system states for each VAR model in the system. We included self-connections in each adjacency matrix to better visualize their generated time series. The number of inter-variable connections between time series was varied across systems, with connections chosen randomly. Our Synthetic Systems were designed to approximate the conditions of the TST dataset (Carlson et al., 2023) discussed in Section 5, with increasing fidelity / complexity (the largest systems had 12 nodes, which corresponded to 12 channels in our preprocessed TST dataset).
Once we defined a system, we sampled recordings and their labels for the dataset. First, a random state vector was drawn from a uniform distribution for each VAR model in the system. This vector was recurrently fed through its corresponding VAR model for a number of ‘burn-in’ steps, after which we recorded the resulting multivariate time series for $T$ time steps; innovations were simulated throughout by multiplying draws from uniform and Gaussian distributions. Recordings obtained from each VAR model were then weighted over time with linearly interpolated weights (randomly selected between 0 and 1). These weighted recordings were added together along with a level of Gaussian noise. We augmented labels by marking with a one-hot vector which factor weight was largest at each time step, meaning all of our Synthetic Systems experiments contain label noise.
@@
where $r_{\mathrm{graph}}$ is the ratio of true inter-variable interactions in each system state to the number of possible inter-variable interactions. In essence, $\mathfrak{C}$ is inversely related to the sparsity of relationships in the graph. We provide additional details regarding this complexity score in Appendices C and D.
@@ 4.2 Hypothesis Generation for Synthetic Systems
Figure 2 gives results from a high-complexity system (6-2-2) and three moderate system configurations for which we generated five random repeats (i.e., unique systems) to train each algorithm (more results in Appendix C). Each repeat was designed to have 1040 training samples and 240 validation samples for each factor/system state in the system, with each sample recording 100 sequential time steps. We used the cMLP and cLSTM (Tank et al., 2021); DGCNN (Song et al., 2018); and dCSFA-NMF (Talbot et al., 2023; Mague et al., 2022) methods as baselines. Hyperparameters for each algorithm were chosen via grid search on a single repeat in our (low-complexity) $\mathrm{n_{c}}=12$-node, $\mathrm{n_{e}}=33$-edge (inter-variable), $\mathrm{n_{k}}=3$-factor Synthetic System dataset. Our grid searches selected parameters which minimized stopping criteria (for DYNOTEARS and REDCLIFF-S) or objective functions (for all other baselines), with ties won by the most ‘parsimonious’ parameters (see Appendices D and E). The same hyperparameters were used across all the experiments reported in this section, with limited adjustments - e.g., to the number of REDCLIFF-S factors or the value of loss coefficients - made to account for certain characteristics in the target system (see Appendix E). Each algorithm was trained on a given system repeat dataset, after which we computed the average optimal f1-score and standard error of the mean (SEM) of each algorithm’s inter-variable causal predictions against the true causal graphs across factors and repeats for each synthetic system.
@@
Table 2: Comparing unweighted inter-variable edge detection by supervised discovery methods on the 12-11-2 Synthetic Systems (Fig. 2) - note REDCLIFF-S’ f1-scores differ from Fig. 2 which included edge weighting. Edge predictions were determined by optimal f1-score thresholds. Values report the mean (over repeats) of the median (over 10 random seeds and system factors) statistic $\pm$ SEM. ‘Upper’ and ‘Lower’ refer to upper- and lower-triangular portions of adjacency matrices, which we split to avoid cycles (see Sec. 4.1 & Appendix B.2). We also report the mean comparative placement (from 1 to 5) of each method’s mean performance (details in Appendix C.3).
@@ 4.3 D4IC Multi-State Extension of DREAM4
Hyperparameters for each algorithm were tuned to a single repeat of the D4IC MSNR dataset prior to training on all D4IC noise levels and repeats. Model selection proceeded as in Section 4.2, although we increased the weight on the cosine similarity penalty in selecting our final REDCLIFF-S model which balanced the scale of variation of the cosine similarity penalty with that of the other penalties (details in Appendix B.3). We include two cMLP baseline models (v1 and v2) with slightly different hyperparameters. Results are visualized in Figure 4, in which we see REDCLIFF-S improves over alternative approaches on this dataset.
@@ 4.4 Discussion
The main strength of REDCLIFF-S is that its hypotheses can be easily tested. Equations 4 and 8 are designed for this in several ways. As discussed, factor scores $\alpha_{n,k,t}$ can be given explicit behavioral designations. Factor scores can also be constrained (via Sigmoid or similar activation functions) to be nonnegative, and thereby be viewed as how (physically) present the corresponding factors are in predicted signals. [^5] Our early experiments included these constraints, but we omit them here due to different models being selected for by our criteria. Thirdly, trained factors $f_{\phi_{k}}$ rely on fixed graphs explicitly showing relationships between variables. Finally, the sum over factors allows us to inspect individual contributions of each factor. Overall, REDCLIFF-S shows performance improvements in many scenarios.
@@
First, Figure 2 and Table 2 together give the core results on our Synthetic Systems Dataset experiments. Figure 2A presents average optimal f1-scores (across all samples and repeats) of REDCLIFF-S and applicable baselines. Surprisingly, the single-factor baselines attain fairly competitive f1-scores in some cases, seemingly learning an ‘average factor graph’ that may not be particularly accurate for any single factor but which performs well ‘on average’. Even so, in Figure 2B we see the average pairwise improvement of REDCLIFF-S’ predictions is at least a standard error of the mean above zero for all baselines on all four systems, suggesting REDCLIFF-S yields improvements with statistical significance. These findings are further supported by the results shown in Table 2. Despite the fact that REDCLIFF-S was the only algorithm trained to predict global state information (all other baselines in Table 2 had direct access to the state label), REDCLIFF-S is uniquely able to achieve competitive performance across all metrics used. This can be seen in that when we rank the algorithms based on their mean performance for each metric, REDCLIFF-S’ ranking (or ‘comparative placement’, to avoid confusion with ‘matrix rank’) tends to be lower than the other methods’. Taken together, our findings in Section 4.2 suggest REDCLIFF-S generates hypotheses about the presence of dynamic causal relationships that prove accurate at state-of-the-art rates, and this pattern holds true for systems with differing numbers of variables, inter-variable relationships, and factors.
@@
Results from our D4IC HSNR experiments are shown in Figure 4, where we see the REDCLIFF-S (cMLP) model achieves the highest average optimal f1-score with a statistically significant margin. While the improvements over competing methods are somewhat modest in this case, a brief investigation into the factor networks of D4IC HSNR suggests that four of the five factors in each repeat would have been rated as ‘low complexity’ according to Equation 11; importantly, REDCLIFF-S tends to make the least improvement over baselines in the ‘low complexity’ regime (see Appendix C.3). In all, our D4IC experimental results generally support those in Section 4.2, verifying that REDCLIFF-S is effective at hypothesizing dynamic causal relationships.
@@ 5 Experiments on Real-World Datasets
We tested REDCLIFF-S on real-world data using the publicly available Tail Suspension Test (TST) dataset (Carlson et al., 2023) - which we discuss here - and the Social Preference (SP) dataset described in Mague et al. (2022) (see Appendix C.5). Where possible, we combined multiple local field potential (LFP) recordings within a single brain region by taking their average. The number of factors and other hyperparameters were chosen based on grid searches minimizing the stopping criteria for training. Appendix D covers TST preprocessing in detail, and Appendix C.4 elaborates on TST model selection and additional results.
In Figure 5 we show the difference between mean normalized causal estimates for the REDCLIFF-S factors assigned to the Open Field (OF) and Home Cage (HC) behavioral paradigms in this case study. The OF factor relies more heavily than the HC factor on information from the Nucleus Accumbens Core (Acb_Core) to predict behavior in the Prelimbic and Infralimbic cortices, and from the Anterior-Left Lateral Habenula (‘alLH_HAB’) and Medial Dorsal Hippocampal (mDHip) channels to predict activity in several regions, including the Thalamus (Md_Thal). Interestingly, our model’s hypothesis that the Anterior-Left Lateral Habenula (‘alLH_HAB’) is important for predicting LFP activity across several brain regions in the more stressful OF paradigm appears to be validated by recent research implicating the Lateral Habenula in stress responses of mice (Tan et al., 2024). Previous research also identified a brain-wide ‘anxiety network’ involving the Nucleus Accumbens Shell and the Prelimbic and Infralimbic cortices (Talbot et al., 2023), which our REDCLIFF-S model appears to express, in part, as pathways from the Nucleus Accumbens Core to the Prefrontal Cortex and from there to the Infralimbic Cortex and then to the Nucleus Accumbens Shell.
@@ 6 Conclusion
We present the novel REDCLIFF-S hypothesis generation algorithm for temporal systems featuring dynamic causal interactions. By learning to combine nonlinear factors using weights conditioned on signal history, REDCLIFF-S attains state-of-the-art performance in estimating multiple causal graphs simultaneously in various settings and in a way conducive to follow-up scientific inquiry. Experiments on real data suggest that REDCLIFF-S can provide scientific value to computational neuroscience and other fields.
@@ Acknowledgements
We would like to thank our colleagues at the Collective for Psychiatric Neuro-Engineering (CPNE) for supporting this work. In particular, conversations with Noah Lanier, Aaron Fleming, Casey Hanks, Kathryn Walder, and Hannah Soliman proved especially helpful in clarifying and communicating our ideas. Conversations with Natalie Brown were also useful for clarifying key ideas. We would also like to thank the reviewers for providing their insight and helpful feedback.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2139754. Research reported in this publication was also supported by the National Institute Of Mental Health of the National Institutes of Health under Award Number R01MH125430. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
@@ A.1 Identifiability Analysis of Systems Featuring Nonlinear, Dynamic Causal Graphs
In this section, we provide theoretical arguments as to why the systems at the center of our work - dynamical systems featuring nonlinear, dynamic causal graphs - are generally not identifiable. At the same time, we point out that prior work has demonstrated that using causal discovery methods can still benefit scientific research in real-world settings where nonlinear, dynamic causal graphs are the norm rather than the exception (Mague et al., 2022). Throughout our discussion, we follow the common definition of identifiability, in which if the likelihood of two models is the same, then their parameterizations must also be the same (Maclaren & Nicholson, 2019).
@@
There are three cases related to causal identifiability in our setting: the case where auxiliary variables are not available (the unsupervised case), that where some - but not all - auxiliary variables are available (the partially supervised case), and the case where an auxiliary variable can be assigned to every system state (the fully supervised case). For the unsupervised and partially supervised cases, the lack of identifiability is readily apparent in that we can swap various factors and factor weights tied to unsupervised states within our REDCLIFF(-S) architecture and arrive at functionally identical models.
@@
The fully supervised case is also not identifiable, as we illustrate via a counterexample. Consider the system with $n_{\mathrm{k}}=2$ states (00 and $1$) and $n_{\mathrm{c}}=2$ nodes ($\mathrm{A}$ and $\mathrm{B}$) where both nodes are self-connected. Assume all innovations in $\mathrm{A}$ and $\mathrm{B}$ follow a symmetric random distribution (e.g., the uniform or normal distributions) in both states 00 and $1$ at each time step. Finally, assume the only difference between states 00 and $1$ is that in state 00 we have that $\mathrm{A}$ drives $\mathrm{B}$ (i.e., $\mathrm{A}\rightarrow\mathrm{B}$) via the ReLU (that is $\mathrm{ReLU}(x)=x\mathbf{1}_{[\mathrm{max}(x,0)]}(x)$) function at lag $t=2$, whereas in state $1$ we have $\mathrm{A}\rightarrow\mathrm{B}$ is the negative ReLU (that is $\mathrm{nReLU}(x)=x\mathbf{1}_{[\mathrm{min}(x,0)]}(x)$) function at lag $t=2$. In this setting, the functional form of the $\mathrm{A}\rightarrow\mathrm{B}$ edge cannot be uniquely identified for either state 00 or $1$ when given an infinite number of samples. This is because for any given sequence of innovations $\mathcal{S}=([\mathrm{A}_{t},\mathrm{B}_{t}])_{t=1}^{T}$, we have that $\exists S^{\prime}=([-\mathrm{A}_{t},\mathrm{B}_{t}])_{t=1}^{T}$ which satisfies
$\mathrm{ReLU}(\mathrm{A}_{t})+\mathrm{B}_{t}=\mathrm{nReLU}(-\mathrm{A}_{t})+\mathrm{B}_{t}$
for any $t\in\mathbb{N}_{\leq T}$. In other words, for every instance in which it appears that $\mathrm{A}$ is influencing $\mathrm{B}$ via the ReLU update rule, we will observe another instance in which $\mathrm{A}$ is possibly influencing $\mathrm{B}$ via the nReLU update with the exact same effect. Thus, the ReLU and nReLU update rules are equally likely in both state 0 and in state 1, rendering the two states unidentifiable by definition.
@@
In summary, we have demonstrated that systems featuring nonlinear, dynamic causal graphs are not identifiable unless one makes sweeping assumptions, which preclude even fairly simple systems (such as our 2-node system with ReLU-like nonlinearities) from being considered. Despite this, previous work has shown that causal discovery methods still benefit scientific research in real-world settings where nonlinear, dynamic causal graphs are common (Mague et al., 2022).